OpenStudy (anonymous):

Fan and medal Factor completely: 2x^2 + 14x + 12 A. (x + 1)(x + 6) B. 2(x + 2)(x + 3) C. 2(x + 1)(x + 6) D. 2(x + 2)(x + 6)

OpenStudy (anonymous):

2x^2 - 14x + 24 divide out common factor (2) x^2 - 7x + 12 Factors of 12 that add to - 7: -3, -4 Factorization: (x-3)(x-4)

OpenStudy (anonymous):

2x6-14x=24 divided out common factor of (2) x^2 -7x +12 factors of 12 that add to -7 -3,-4 so the it will be (x-3)(x-4)

OpenStudy (anonymous):

i mean ^2

OpenStudy (anonymous):

Thanks for repeating exactly what he said

OpenStudy (anonymous):

2(x + 2)(x + 3)

OpenStudy (anonymous):

Answer is C

OpenStudy (anonymous):

2(x+1)(x+6)

OpenStudy (anonymous):

2(x^2+ 6x + 1x + 6) =2(x+1)(x+6)